OpenStudy (anonymous):

how much energy is needed to raise the temperature of a silver necklace chain with a mass of 22.5 g from room temperature, 25 degrees celsius to a body temperature of 37 degrees celsius

OpenStudy (accessdenied):

We need to know one piece of information: the specific heat of Silver metal. This is the amount of energy to raise the temperature of 1 gram of material by 1 Celsius degree. \( \dfrac{\text{J}}{\text{g $\cdot$ C$ ^o$}} \) Notice that by observing these units, we simply have to multiply specific heat by the change in temperature and the mass: \( \text{Energy} = m \times \Delta T \times C \) The end unit will be in Joules, the unit of energy!